OpenStudy (anonymous):

Use vertices and asymptotes to graph the hyperbola. Find the equations of the asymptotes. y = ± square root of x^2 - 5

OpenStudy (anonymous):

A. Asymptotes: y = ± x B. Asymptotes: y = ± 5/3 x C. Asymptotes: y = ± 5/3 x D. Asymptotes: y = ± x

OpenStudy (anonymous):

so it's \(y^2=x^2 - 5\), \(x^2- y^2= 5\) asymptotes are y = ± x and it's a horizontal hyperbola.

OpenStudy (anonymous):

so it would be d?

OpenStudy (anonymous):

yup.

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

you're welcome :)

OpenStudy (anonymous):

It was incorrect.

OpenStudy (anonymous):

this would help. http://www.wolframalpha.com/input/?i=y%5E2%20%3D%20%20x%5E2%20-%205&t=crmtb01

OpenStudy (anonymous):

the answer is d.